OpenStudy (anonymous):

solve using substitution y=2x+9 y=x+6

OpenStudy (a1234):

y=2x+9 y=x+6 Write one equation like this: -2x + y = 9 -2x + x + 6 = 9 -x = 3 or x = -3 y = -3 + 6 y = 3 (-3, 3)